OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8

hero (hero):

4x + y = 11 x + 2y = 8 y = 11-4x y = (8-x)/2 y = y (substitute appropriate expressions) 2(11-4x) = 8-x 22 - 8x = 8-x 14 = 7x x = 2 y = 3